OpenStudy (amurciano2):

A person standing at the edge of a seaside cliff kicks a stone over the edge with a speed of 20 m/s. The cliff is 54 m above the water’s surface, as shown.With what speed does the stone strike the water? Answer in units of m/s

OpenStudy (anonymous):

 As you can see from the drawing the Vox is 20m/s. So by knowing that Vox=20m/s you know that your Vfx = 20m/s because it is constant. Therefore you already know the X component of the final velocity at which it strikes the water. Now you need to find the Y component. \[Vy ^{2} = V _{oy}^{2} - 2g(y-y _{o})\] Now you replace your known values to get your \[Vy\] \[Vy ^{2} = 0 - 2(9.81m/s^{2})(-54m)\] \[\sqrt{1059.48} =32m/s\] Now you know the Vy component. Now to get the magnitude you use the formula \[\sqrt{A _{x ^{2}}+{A _{y ^{2}}}}\] \[\sqrt{(20 m/s){ ^{2}}+{(32.5{m/s) ^{2}}}}\] =38.16m/s That is the velocity at which it strikes the water. If you have any questions I'll be here to answer.